Greetings~
I need to evaluate the variable expression when a= -2, b=4, c= -1, and d=3.
on problem -2/3d- 1/5(bd-ac)=
This is what I have tried:
Solution
distribute the values to the variables:
-2/3(3) - 1/5(4(3) - (-2)(-1))=
simplify equation with parentheses:
-2/3(3) - 1/5(12 -3)=
simplify more:
-2/3(3) - 1/5(9)=
simplify multiplication:
-2/3 * 3/1 + -1/5 * 9/1=
multiply numerators and denominators:
(-2 * 3)/(3 * 1) + (-1 * 9)/ (5 * 1)=
simply:
-6/3 + -9/5=
find common denominator:
(-6*5)/(3*5) + (-9*3)/(5*1)=
simplify:
-30/15 + -27/15=
add numerators:
(-30 + -27)/15=
simplify:
-57/15=
find lowest common multiple:
(-57/3)/(15/3)=
simplify:
-19/5
Thanks =)

Found 2 solutions by stanbon, rfer:
Answer by stanbon(48545)

evaluate the variable expression when a= -2, b=4, c= -1, and d=3.
on problem (-2/3)d- (1/5)(bd-ac)
---
= (-2/3)(3) - (1/5)(4*3-(-2)(-1))
---
= -2 - (1/5)(12 - 2)
---
= -2 -(1/5)(10)
---
= -2-2
---
= -4
=============
Cheers,
Stan H.

Answer by rfer(10417)

looks good
Bob